OpenStudy (anonymous):

Which of the following represents inversevariation? Y = 9/x Y = 3 + x Y/x = -10 Y = 1/6x

OpenStudy (mathstudent55):

Direct variation: y = kx Inverse variation: y = k/x where k is a number

OpenStudy (anonymous):

So A? @mathstudent55

OpenStudy (anonymous):

Help @mathstudent55

OpenStudy (mathstudent55):

Correct. A is it!

OpenStudy (mathstudent55):

I have a question on choice D. Is it \(y = \dfrac{1}{6x}\) or \(y = \dfrac{1}{6}x\) The first one is inverse variation. The second one is direct variation.

OpenStudy (anonymous):

Thanks you so much!

OpenStudy (mathstudent55):

You're welcome.